[figure: 9.7 cm]

STABILITY OF STRIPE PHASES IN THE HUBBARD MODEL

Dariusz Góra, Krzysztof Rościszewski, and Andrzej M. Oleś Institute of Physics, Jagellonian University, Reymonta 4, PL-30059 Kraków, Poland

Motivated by the experimental discovery of stripe phases in the cuprates [1], we investigate the stability and magnetic ordering in various stripe phases obtained for the two-dimensional Hubbard model with and without static phonons using finite 8 × 8 clusters with periodic bondary conditions solved within the Hartree-Fock (HF) approximation and in a variational local ansatz. The magnetic order is antiferromagnetic (AF) at half-filling (n = 1), and it changes to a nonhomogeneous charge and magnetization distributions at finite hole doping δ > 0. Nonmagnetic vertical (01) stripes, found before as only locally stable in the HF approximation [2], are stabilized by local correlations and give the ground states for the underdoped (δ = 1/8) and overdoped (δ = 1/4) systems. At both doping levels the charge unit cell is by a factor of two smaller than the magnetic one and consists of four atoms, with domain walls separating AF domains and filled by one hole per two (one) wall atoms at δ = 1/8 (δ = 1/4). Finite next neighbor hopping elements, t′ and t′′, stabilize extended diagonal stripes, similar to those found in the dynamical mean-field theory [3]. These findings show that several types of stripe ordering compete with each other in the regime relevant for the cuprates in qualitative agreement with the experimental results [1].

[1] J. M. Tranquada et al., Phys. Rev. B 54, 7489 (1996).

[2] J. Zaanen and A. M. Oleś, Ann. Phys. (Leipzig) 5, 224 (1996).

[3] A. I. Lichtenstein et al., J. Supercond. 12, in press (1999).

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : poster

Corresponding author : Dariusz Góra

Address for correspondence : Institute of Physics, Jagellonian University, Reymonta 4, PL-30059 Kraków, Poland

Email address : dgora@jetta.if.uj.edu.pl